The area of a rectangle is 54 cm the length is 2 cm more than a X and the width is 5 cm less than twice X solve for X run your answer to the nearest whole number
Answer by CubeyThePenguin(3113)

(x + 2)(2x - 5) = 54
2x^2 - x - 10 = 54
2x^2 - x - 64 = 0


Use the quadratic formula: x = -5.4124 or 5.9124.

Distance can't be negative, so x is roughly 6 cm.